SUPERFLUID EFFECTS IN PARA-H_{2} CLUSTERS PROBED BY CO_{2} ROTATION-VIBRATION TRANSI-TIONS

HUI LI, ROBERT J. LE ROY, PIERRE-NICHOLAS ROY, Department of Chemistry, University of Waterloo, Waterloo, ON N2L 3G1, Canada; \underline{A.R.W. MCKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

The prospect of directly observing superfluidity in para-H_{2} is a tantalizing but elusive goal. Like ^{4}He, para-H_{2} is a light zero-spin boson. However, H_{2}-H_{2} intermolecular interactions, though weak, are stronger than He-He interactions, and hydrogen is a solid below about 14 K. This makes detection of superfluidity in bulk hydrogen problematical, to say the least. But there are still possibilities for para-H_{2} in the form of clusters or in nano-confined environments, and superfluid transition temperatures as high as ∼6 K have been predicted.^{a} Spectroscopic observations of (para-H_{2})_{N}-CO_{2} clusters^{b} were at first very difficult to interpret for N > 5. However, with the help of path integral Monte Carlo simulations and an accurate new H_{2}-CO_{2} intermolecular potential surface^{c} which explicitly incorporates dependence on the CO_{2} ν_{3} asymmetric stretch, it is now possible to achieve a remarkably consistent picture of (para-H_{2})_{N}-CO_{2} clusters in the size range N = 1 ∼ 20. By combining the experimental spectroscopic measurements and theoretical simulations, we determine the size evolution of the superfluid response of the CO_{2}-doped para-H_{2} clusters, which peaks for the “magic” number N = 12.

aV.L. Ginzburg and A.A. Sobyanin, JETP Lett. 15, 343 (1972).

bA.R.W. McKellar, Paper WH04, 63rd OSU International Symposium on Molecular Spectroscopy, June 16-20, 2008.

cH. Li, P.-N. Roy, and R.J. Le Roy, J. Chem. Phys., submitted.